OpenStudy (anonymous):

Find formulas for the following inverse functions. a. Let f(x) = 8 -2x Find f^-1(y) b. g(t) = sqrt(2-t) Find g^-1(s) c. f(t) = 1000e^0.2t Find f^-1(p)

OpenStudy (anonymous):

I had similar problems on my exam and have no clue how to solve.

OpenStudy (anonymous):

how would you solve \[8-2x=7\]?

OpenStudy (anonymous):

I remember that inverse is the opposite of y so I guess you are looking for x

OpenStudy (anonymous):

I would subtract 8 from both sides and get -2x = -1 and then divide by -2

OpenStudy (anonymous):

The word inverse is what I don't understand

OpenStudy (anonymous):

right a) subtract \(8\) b) divide by \(-2\)

OpenStudy (anonymous):

your inverse is therefore \[f^{-1}(y)=\frac{y-8}{-2}\]

OpenStudy (anonymous):

or if you prefer \[f^{-1}(y)=\frac{8-y}{2}\] same thing

OpenStudy (anonymous):

so is the problem just asking me to solve like I normally would do?

OpenStudy (anonymous):

you can find it in your head, or you can also write \[8-2x=y\] and solve for \(x\) via \[8-2x=y\\ y-8=-2x\\ x=\frac{y-8}{-2}\\ f^{-1}(y)=\frac{y-8}{-2}\]

OpenStudy (anonymous):

For b, I would have sqrt(2) = y + t

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

\[ g(t) = \sqrt{2-t} \]

OpenStudy (anonymous):

how would you solve \[\sqrt{2-t}=5\]?

OpenStudy (anonymous):

you would square both sides

OpenStudy (anonymous):

yes giving \(2-t=5^2\) then what?

OpenStudy (anonymous):

@satellite73 , is that right?

OpenStudy (anonymous):

f^-1(y) = (2-t)^2

OpenStudy (anonymous):

@sourwing

OpenStudy (anonymous):

your answer above to the question is wrong you are not done solving for \(t\) yet in \(\sqrt{2-t}=5\) first step was right, square both sides to get \[2-t=5^2\]

OpenStudy (anonymous):

then subtract \(2\) and get \[-t=5^2-2\] then change the sign and get \[t=2-5^2\]

OpenStudy (anonymous):

now we can solve \[s=\sqrt{2-t}\]for \(t\) square both sides \[s^2=2-t\] subtract \(2\) \[s^2-2=-t\] change the sign \[2-s^2=t\] and your inverse is \[g^{-1}(s)=2-s^2\]

OpenStudy (anonymous):

(x)^2 = (sqrt(2-y))^2 x^2 = 2 - y x^2 - 2 = -y

OpenStudy (anonymous):

@satellite73